OpenStudy (anonymous):

how do you do y^2 - 28 = 0 using the square root property of equality

hero (hero):

Square Root Property of Equality: If y^2 = a, then sqrt{y^2) = sqrt{a}. On the left side, the sqrt and square are inverses of each other so they cancel. On the right side, sqrt{a} = + or - sqrt{a} For example, if y^2 = 4, then sqrt{y^2} = sqrt{4}, which becomes y = + or - 2

OpenStudy (anonymous):

where does the "a" come from?

hero (hero):

It's a general case

hero (hero):

a means any number